i have this problem
2e^(4x-2)=9
i know that i can use the calculator to find (e) but what about x?
thank you in advance! :-]

Answer by nerdybill(5413)

You must apply "log rules":
2e^(4x-2)=9
e^(4x-2) = 9/2
Take the natural ln of both sides:
4x - 2 = ln(9/2)
4x = ln(9/2) + 2
x = [ln(9/2) + 2]/4
x = 0.876